OpenStudy (anonymous):

Which real number lies farther from -4 on the real number line 3 or -10? Explain.

OpenStudy (anonymous):

A couple ways you can do this. you can draw a number line and count or you can do -4+x=3 and -4+x=-10 and find x

OpenStudy (anonymous):

If you go from -4 to 10, Then thats 6 Numbers. but if you go from -4 to 3, Thats 7 numbers. So which is farther?

OpenStudy (anonymous):

3 @wrstlr3232

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

Yup! :3

OpenStudy (anonymous):

If you go from -4 to 10, Then thats 6 Numbers @MarrBear

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

How can i explain it ? -4+x=3 and -4+x=-10 is 3. is that good @MarrBear @wrstlr3232

OpenStudy (anonymous):

Remember that your answer for x is going to be the absolute value. So if you find -4+x=-10 you'll have -6. But it's really 6 away. You're just going in the negative direction.

OpenStudy (anonymous):

How can i explain it ? -4+x=3 and -4+x=-10 is 3. is that good @wrstlr3232

OpenStudy (anonymous):

Yeah that works. You can solve for x if you want. It depends on what your teacher would want. I would say -4+x=3 and -4+x=-10 and then x=7 and x=-6. So 3 is further away because the x value is a larger number.